OpenStudy (anonymous):

Simplify f(x) = (3x^2-4x+1)/(3x-1). Using complete sentences, explain why f(1) = 0, f(0) = -1, and f(-1) = -2, yet f(1/3) is undefined. please show how u do it so i can understand

OpenStudy (anonymous):

(3x^2-4x+1)/(3x-1). lets try to factor numerator (3x 1)(x 1) to negative will work (3x-1)(x-1)

OpenStudy (anonymous):

(3x-1)(x-1) -------------- 3x-1 simplifies to (x-1)

OpenStudy (turingtest):

To see why f(1) = 0, f(0) = -1, and f(-1) = -2 just plug in the numbers into the function i.e. \[f(x)=\frac{3x^2-4x+1}{3x-1}\]with x=1 becomes\[f(1)=\frac{3(1)^2-4(1)+1}{3(1)-1}=\frac{0}{2}=0\]do the same with x=0 and x=-1 and you will get the expected value for f(x) Now try it for f(1/3):\[f(1/3)=\frac{3(1/3)^2-4(1/3)+1}{3(1/3)-1}=\frac{0}{0}\]which is undefined, because anything divided by 0 is undefined.

OpenStudy (turingtest):

The denominator is where to look to find where a function is undefined. Note that imranmeah's factorization of the function is not valid at the point x=1/3, so even though it can be simplified for practical purposes, you must always consult the original function to see where it may be undefined.

OpenStudy (anonymous):

thank you so much. I think I actually understand how to do it now.